# Optimal Neural Compressors for the Rate-Distortion-Perception Tradeoff

### v1

@@ Appendix C Proofs
$=\frac{1}{2(\sigma^{2}-\nu)}\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{y}\|^{2}\right]+\frac{1}{n}\log\rho_{\sqrt{\sigma^{2}-\nu}}(\Lambda)+\epsilon^{\prime\prime}$ (162)
$=-\frac{1}{n}\sum_{\bm{\lambda}\in\Lambda}q_{\bm{y}}(\bm{\lambda})\log q_{\bm{y}}(\bm{\lambda})+\epsilon^{\prime\prime}$ (163)
$=\frac{1}{n}H(\bm{y})+\epsilon^{\prime\prime},$ (164)
for $n$ sufficiently large, where (161) is by (144) and (158). Therefore, for $n$ sufficiently large, we have
$\frac{1}{n}H(Q(\alpha X^{n}))$ $=\frac{1}{1-\epsilon^{\prime\prime}}\left[\frac{1}{n}H(\bm{y})+\epsilon^{\prime\prime}\right]$ (165)
$\leq\frac{1}{1-\epsilon^{\prime\prime}}\left[\frac{1}{2}\log\frac{\sigma^{2}}{\nu(1+\epsilon_{2})}+\epsilon^{\prime}+\epsilon^{\prime\prime}\right]$ (166)
$\leq\frac{1}{1-\epsilon^{\prime\prime}}\left[\frac{1}{2}\log\frac{\sigma^{2}}{\nu}+\epsilon^{\prime}+\epsilon^{\prime\prime}\right]$ (167)
$=\frac{1}{1-\epsilon^{\prime\prime}}\left[\frac{1}{2}\log\frac{D^{2}-4\sigma^{2}D+8\sigma^{4}}{D(4\sigma^{2}-D)}+\epsilon^{\prime}+\epsilon^{\prime\prime}\right]$ (168)
where (165) is due to (135) and (164), and (166) holds by Ling and Belfiore [2014, Lemma 6] for any $\epsilon^{\prime}>0$ and $n$ sufficiently large. Thus the result follows by taking $\epsilon^{\prime},\epsilon^{\prime\prime}\rightarrow 0$ with $n\rightarrow\infty$.
Appendix D Empirical Evaluation of Sliced Wasserstein

### v2

@@ Appendix D Proofs
$=\frac{1}{2(\sigma^{2}-\nu)}\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{y}\|^{2}\right]+\frac{1}{n}\log\rho_{\sqrt{\sigma^{2}-\nu}}(\Lambda)+\epsilon$ (178)
$=-\frac{1}{n}\sum_{\bm{\lambda}\in\Lambda}q_{\bm{y}}(\bm{\lambda})\log q_{\bm{y}}(\bm{\lambda})+\epsilon$ (179)
$=\frac{1}{n}H(\bm{y})+\epsilon,$ (180)
for $n$ sufficiently large, where (177) is by Lemma D.4. Therefore, for $n$ sufficiently large, we have
$\frac{1}{n}H(Q(\alpha X^{n}))$ $=\frac{1}{1-4\epsilon_{z}}\left[\frac{1}{n}H(\bm{y})+\epsilon\right]$ (181)
$\leq\frac{1}{1-4\epsilon_{z}}\left[\frac{1}{2}\log\frac{\sigma^{2}}{\nu(1+\epsilon_{2})}+\epsilon_{h}+\epsilon\right]$ (182)
$\leq\frac{1}{1-4\epsilon_{z}}\left[\frac{1}{2}\log\frac{\sigma^{2}}{\nu}+\epsilon_{h}+\epsilon\right]$ (183)
where (181) is due to (170) and (180). (182) holds by Ling and Belfiore [2014, Lemma 6], the choice of volume, for $n$ sufficiently large.
We now prove Thm. 4.5, which is restated below.
Theorem D.6 (Optimality of PD-LTC for Gaussian sources (Thm. 4.5 in main text)).
$\lim_{n\rightarrow\infty}\frac{1}{n}H(Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n})))\leq R\left(\nicefrac{{D}}{{2}},\infty\right),$ (184)
$\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}[\|X^{n}-\hat{X}^{n}\|^{2}]\leq D,$ (185)
$\lim_{n\rightarrow\infty}\frac{1}{n}W_{2}^{2}(P_{X^{n}},P_{\hat{X}^{n}})=0,$ (186)
where $\hat{X}^{n}=g_{s}^{(n)}(Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n}))+s\bm{u})$, $\bm{u}\sim\mathrm{Unif}(\mathcal{V}(\Lambda^{(n)}))$, and $s=\frac{\sigma}{\sqrt{\sigma^{2}-\nicefrac{{D}}{{2}}}}$.
Proof.
Set $s=\frac{\sigma}{\sqrt{\sigma^{2}-\nu}}$, $\beta=1$, and $\nu=\frac{D}{2}$, which satisfies (88). Choose the lattices $\Lambda^{(n)}$ as polar lattices [Liu et al., 2021, 2024]. By the choice of $s$, $\epsilon_{u}=\epsilon_{z}=\epsilon_{\Lambda^{(n)}}\left(\sqrt{\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}}\right)$, and therefore both flatness factors will vanish exponentially fast as $n\rightarrow\infty$ by Liu et al. [2021, Prop. 1]. Additionally, by using the fact that $\epsilon_{1}\leq\epsilon_{\Lambda^{(n)}}\left(\nicefrac{{\sqrt{\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}}}}{{\sqrt{\frac{\pi}{\pi-t}}}}\right)$, and since $\epsilon_{z}$ vanishes, taking $t\rightarrow 0$ results in a vanishing $\epsilon_{1}$ and therefore vanishing $\epsilon_{h}$; see Ling and Belfiore [2014, Sec. III]. By Liu et al. [2019], polar lattices are AWGN-good and have exponentially decaying MAP error probability for the $\tilde{\bm{x}}=\bm{y}+\bm{z}$ AWGN channel with lattice Gaussian input, at any signal-to-noise ratio (and therefore any $\nu\in(0,\sigma^{2})$). By Liu et al. [2024], the polar lattices are also quantization-good. Therefore, by Lemma D.5 we have that the rate, distortion, and perception satisfy
$\lim_{n\rightarrow\infty}\frac{1}{n}H(Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n})))\leq\frac{1}{2}\log\frac{2\sigma^{2}}{D},$ (187)
$\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}[\|X^{n}-\hat{X}^{n}\|^{2}]\leq D,$ (188)
$\lim_{n\rightarrow\infty}\frac{1}{n}W_{2}^{2}(P_{X^{n}},P_{\hat{X}^{n}})=0,$ (189)
when the transforms are chosen to be $g_{a}^{(n)}(\bm{v})=\alpha\bm{v}$ and $g_{s}^{(n)}(\bm{v})=\beta\bm{v}$. ∎
Remark D.7.
In the proof of Lemma D.5 (and therefore Thm. 4.5), the condition in (88) ensures that the distribution of $\hat{X}^{n}$ is approximately Gaussian with covariance $\sigma^{2}I_{n}$, and the closeness of this approximation is quantified by $\epsilon_{u}$. This essentially controls how well the perception constraint can be enforced, which is why $s$ shows up in the flatness factor. One can see that $s$ needs to be sufficiently large to ensure $\epsilon_{u}=\epsilon_{z}$ which guarantees $\epsilon_{u}$ vanishes by Liu et al. [2021, Prop. 1]. If $s$ is too small, $\epsilon_{u}>\epsilon_{z}$ and it is not guaranteed it will vanish. We see that the choice of $s=\frac{\sigma}{\sqrt{\sigma^{2}-\nicefrac{{D}}{{}}2}}>1$ indicates a dither $s\bm{u}$ that “bleeds” outside the lattice cell at low rates is sufficient to ensure a vanishing $\epsilon_{u}$ and therefore vanishing perception. This seems to imply it is not possible to choose a smaller $s$ than the one chosen and still make $\epsilon_{u}$ vanish; if this were possible, it would imply that a near-perfect perception rate-distortion tradeoff below $R(\nicefrac{{D}}{{2}},\infty)$ is achievable, which would contradict results in the RDP literature that $R(\nicefrac{{D}}{{2}},\infty)$ is the lower bound [Yan et al., 2021, Wagner, 2022, Chen et al., 2022].
On the other hand, $\epsilon_{z}$ is used to approximate $P_{X^{n}}\approx P_{\tilde{\bm{x}}}$, and $\epsilon_{h}$ is used to approximate the entropy rate of $\bm{y}$; as shown in the proof, a vanishing $\epsilon_{z}$ implies a vanishing $\epsilon_{h}$.
Remark D.8.
The sequence of PD-LTCs in Thm. 4.5 that satisfy the near-perfect perception constraint in (16) can be upgraded to a sequence of codes satisfying a perfect perception constraint of $P_{X^{n}}=P_{\hat{X}^{n}}$ with the same asymptotic rate and distortion, by following the “coupling argument” of Saldi et al. [2015]. That is, if $\pi_{X^{n},\hat{X}^{n}}$ is the coupling that satisfies $\frac{1}{n}W_{2}^{2}(P_{X^{n}},P_{\hat{X}^{n}})=\epsilon$, then one can use $\bar{X}^{n}\sim\pi_{X^{n}|\hat{X}^{n}}$ as the reconstruction instead of $\hat{X}^{n}$. This would ensure $P_{\bar{X}^{n}}=P_{X^{n}}$ and the distortion $\frac{1}{n}\mathbb{E}[\|X^{n}-\bar{X}^{n}\|^{2}]\leq\frac{1}{n}\mathbb{E}[\|X^{n}-\hat{X}^{n}\|^{2}]+\epsilon$. Since this change only affects the decoder, the rate remains the same as before.
Remark D.9.
The choice of polar lattices in the proof of Thm. 4.5 may not be necessary, and other lattice families may also work. The essential requirements are that both $\epsilon_{u}$ and $\epsilon_{z}$ vanish, AWGN-goodness, quantization-goodness, and vanishing MAP error probability for the $\tilde{\bm{x}}=\bm{y}+\bm{z}$ AWGN channel. While the latter three can be simultaneously satisfied by other lattice families, such as the mod-$p$ lattices [Loeliger, 1997], to the best of the authors’ knowledge, the only currently known lattice family with known results on vanishing $\epsilon_{u}$ and $\epsilon_{z}$ is the polar lattice. This does not preclude the existence of other lattice families that may satisfy all the aforementioned requirements.
Appendix E Empirical Evaluation of Sliced Wasserstein